How many grams of CO2 would be made from 3,752 g Ti(HCO3)4 of if the reaction proceeds to 93%?
Ti(HCO3)4 + HBr  TiBr4 + 4 [ H2CO3 ]  4 H2O + 4 CO2
1
Expert's answer
2017-11-13T09:00:06-0500
Ti(HCO3)4 + 4 HBr = TiBr4 + 4 H2O + 4 СО2
3.752 g x g
292 g (4x44) g

X = (3.752 x 4x44) / 292 = 2.261 g ‒ m(theoretical)
eta = m(practice) / m(theoretical) = 0.93
m(practice) = m(theoretical) x eta = 2.261 x 0.93 = 2.103 g